# Structure Language Models for Protein Conformation Generation

### v1

@@ 4.2 Conditional Masked Diffusion Language Model
For conformation generation, we now consider the conditional case of masked diffusion. Given the amino acid types ${\bm{c}}$, our goal is to sample structure tokens through Eq. 6, utilizing a conditional posterior $q({\bm{z}}_{s}|{\bm{z}}_{t},{\bm{z}}_{0};{\bm{c}})$. This posterior can be re-parameterized similarly by incorporating the condition into the backbone model, resulting in $p_{\theta}({\bm{z}}_{s}|{\bm{z}}_{t};{\bm{c}})=q({\bm{z}}_{s}|{\bm{z}}_{t},{\bm{u}}_{\theta}(t,{\bm{z}}_{t},{\bm{c}}))$. To achieve this goal, the reverse process simulated by $p_{\theta}({\bm{z}}_{s}|{\bm{z}}_{t};{\bm{c}})$ must effectively approximate the data distribution $p({\bm{z}}|{\bm{c}})$. A feasible training objective is to optimize the estimation of the conditional ELBO within the continuous-time integral, resulting in the following loss (see Appendix LABEL:app:loss for details):
@@ 4.3 Bidirectional Encoder as denoising network
We now discuss the implementation of the conditional denoising network using bidirectional encoder language models, such as BERT (Devlin, 2018). First, consider the sequential generalization of masked diffusion with a sequence of categorical variables. Let ${\bm{z}}_{t}$ now be a sequence of discrete structure tokens $\left[{\bm{z}}_{t,[1]},{\bm{z}}_{t,[2]},\dots,{\bm{z}}_{t,[L]}\right]$ where ${\bm{z}}_{t,[i]}\in\bar{V},\forall i=1,\dots,L$. Due to the interpolation scheme of Eq. 4 and Eq. 6, we assume conditional independence and factorize the posterior distribution $p_{\theta}({\bm{z}}_{s}|{\bm{z}}_{t};{\bm{c}})$ across the $L$ output tokens, such that $p_{\theta}({\bm{z}}_{s}|{\bm{z}}_{t},{\bm{c}})=\prod_{i=1}^{L}p_{\theta}({\bm{z}}_{s,[i]}|{\bm{z}}_{t},{\bm{c}})=\prod_{i=1}^{L}q\left({\bm{z}}_{s,[i]}|{\bm{z}}_{t,[i]},{\bm{u}}_{\theta,[i]}(t,{\bm{z}}_{t},{\bm{c}})\right)$ where ${\bm{u}}_{\theta,[i]}(t,{\bm{z}}_{t},{\bm{c}})$ represents the $i$-th output channel of neural network. This implement coincides with the BERT-style transformer architecture and allow us to take advantage of existing protein foundation model, for example ESM3 (Hayes et al., 2024). For a sequence of tokens, the masked $\log$-term in the training objective from Eq. 7 is replaced by the summation: $\sum_{i=1}^{L}\lambda_{M}({\bm{z}}_{t,[i]})\log\langle{\bm{u}}_{\theta,[i]}(t,{\bm{z}}_{t},{\bm{c}}),{\bm{z}}_{0,[i]}\rangle$, with notations the same as defined above.
@@ 5 Experiments
Base settings. We start with the pre-trained dVAE established in Hayes et al. (2024) as the structure tokenizer (frozen). The structure quantization is perform residue-level with a receptive field over local geometric neighborhoods of protein structure. The structure language models as described in Section 3.2 are based off state-of-the-art language models as follows: (1) S-T5 (384M) adopts the T5 (Raffel et al., 2020) architecture with a bidirectional encoder and an autoregressive decoder. (2) S-GPT (961M) models the joint distribution with an uni-directional decoder like GPT2 (Radford et al., 2019). (3) ESM3 (zero shot) is a pre-trained BERT model over multi-modality data and perform zero-shot inference by Gibbs sampling. (4) ESMDiff is a fine-tuned variant of (3) on the PDB data using the masked diffusion objective in Eq. 7. For S-T5 and S-GPT, we embed the sequence tokens with the pre-trained ESM3-1.4B encoder to provide model condition. For ESMDiff, two different sampling paradigm: Gibbs and DDPM are considered (see Appendix B for details).
Table 1: Evaluation results of different conformation generation methods on generating the BPTI conformations. JS-* represents the Jensen-Shannon (JS) divergence between the sampled ensemble and long MD ensemble on pairwise distance (PwD), time-lagged independent components (TIC), and radius of gyration (Rg). The validity as the frequency of clash-free samples in the ensemble is also included. Moreover, the ensemble TM-score and ensemble RMSD are reported with respect to the five kinetic clusters in Shaw et al. (2010). The best results among the SLM family are bold.
@@
Baselines. We consider multiple open-source models as evaluation baselines for the protein multiple conformation generation, which are mainly categorized into (1) MSA-based methods that includes MSA-Subsampling (Del Alamo et al., 2022; Bryant & Noé, 2024) and AlphaFlow (Jing et al., 2024). These methods rely on inference-time retrieval of multiple sequence alignments (MSA) ; (2) Single sequence-based methods: EigenFold (Jing et al., 2023) leverages a harmonic diffusion process conditioned on OmegaFold (Wu et al., 2022) embeddings to generate protein structures, Str2Str (Lu et al., 2024) simulates a round-trip local diffusion conditioned on input structure to explore hypothetical conformations, and ESMFlow (Jing et al., 2024) replaces AlphaFlow with ESMFold as the backbone; (3) Specially tailored for intrinsically disordered proteins (IDPs) generation of idpGAN (Janson et al., 2023). Results reported for baselines are obtained by re-running the inference pipeline and based on their open-source codes. The detailed pipeline can be found in Appendix A.
Training data. The training data for structure language models are controlled to contain only PDB entries on or before May 1st, 2020. This cutoff is aligned with previous works (Jing et al., 2023; 2024; Lu et al., 2024; Hayes et al., 2024) trained on PDB data to make fair comparison. The training set is further filtered to include all monomeric structures with a max resolution of $5.0\mathrm{\AA}$, length ranging from 10 to 1000, which forms a total size of $|{\mathcal{D}}|=112.4\textrm{k}$ as the training data.
@@ 5.1 Structural dynamics of BPTI
In the first experiment, we evaluate the models by generating the conformations of protein bovine pancreatic trypsin inhibitor (BPTI). The structural dynamic patterns of BPTI are well acknowledged in Shaw et al. (2010) with 1ms-long MD simulations, based on which five kinetic clusters have been revealed. Similar to Lu et al. (2024), we report the Jensen Shannon (JS) divergence for distributions of pairwise distance (PwD), time-lagged independent components (TIC), and radius of gyration (Rg); the clash-free validity and the ensemble TM-score and root-mean-square-deviation (RMSD) w.r.t. the kinetic clusters. The benchmark results are shown in Table 5. Following Wang et al. (2024), we also evaluate the best distance of the generated samples to each cluster, as shown in Table 5.1. The RMSD and TM-score are both calculated using the TM-score binary (Zhang & Skolnick, 2004) with structural alignment. Note that the Cluster 3 is a difficult remote folding mode (Wang et al., 2024), yet SLMs achieve a significant improvement by modeling with a smaller matching RMSD.
@@
5.2 Conformation changing pairs
We continue the study on the task of modeling and predicting conformational changes in structural proteins. The authors of Jing et al. (2023) have curated two benchmarking sets of pairing data including (1) 77 pairs of fold-switching proteins (Chakravarty & Porter, 2022); and (2) 90 apo/holo pairs with ligand-induced conformational change (Saldaño et al., 2022) to evaluate the modeling capacity of conformation diversity. Following the setting and evaluation metrics of Jing et al. (2023), we randomly sample a few-shot ensemble with five structures per target and evaluate them based on the correlation metrics of residue flexibility and the ensemble TM-score (Zhang & Skolnick, 2004). The evaluation results for both test sets are shown in Table 5.2.
@@ 5.2 Conformation changing pairs
5.3 Intrinsically disordered proteins
@@ 5.4 Runtime analysis
Figure 5: Runtime profiling for SLMs and baseline methods.
@@ 6 Conclusion and Limitations
In this work, we propose a novel conformation generation framework of structure language models (SLMs). The overall task is divided into two stages: conditional sampling of latent structure tokens and roto-translation invariant structure decoding. We develop and train a variety of conditional language models, especially the masked diffusion-based ESMDiff, which is a fine-tuned variant to enhance the capabilities of ESM3 adapting for conformation generation. Unlike existing methods, SLMs perform amortized distribution learning within an invariant latent space, leading to more efficient inference. By alleviating the need for geometric modeling, SLMs can fully exploit the scalability of modern language model architectures and take advantage of advanced hardware optimizations. Benchmarking results across various conformation generation tasks demonstrate the compelling performance and application potential of SLMs. In summary, the proposed method opens up an intriguing and novel research direction for related communities to explore.
Limitations. The current study presents several limitations worth exploring for future works. Firstly, one can design more advanced dVAE architecture to balance between structure disentanglement and reconstruction fidelity. In addition to the discrete latent space, continuous latent space can also be considered, for example using the latent diffusion models (Rombach et al., 2022). Secondly, it is worthwhile exploring alternative SLM instances specially tailored for the sequence-to-structure translation in consideration of proper inductive biases. Lastly, the outstanding performance of MSA-based methods also indicates the potential to build MSA-conditioned structure language models.
@@ A.1 Baseline evaluations
For baselines, we compare our structure language models against: EigenFold (Jing et al., 2023), Str2Str (Lu et al., 2024), MSA subsampling (Del Alamo et al., 2022), AlphaFlow & ESMFlow (Jing et al., 2024), and idpGAN (Janson et al., 2023). For MSA subsampling, we leverage the official repository of Del Alamo et al. (2022) under AlphaFold v2.3.2 with an MSA depth of 32, i.e. 32 maximum extra MSAs and 16 maximum MSA clusters (half the depth value (Del Alamo et al., 2022)). For AlphaFlow and ESMFlow, we use the base PDB model (no distillation) with no resampling. The input MSAs for test proteins are queried via the ColabFold server (Mirdita et al., 2022). All other parameters are set to the recommended values according to the original implementation by the authors (Jing et al., 2024). For EigenFold, we follow the pipeline in Jing et al. (2023) by using OmegaFold (Wu et al., 2022) to make embeddings for the test set and sampling via the pre-trained diffusion model with default parameters, where $\alpha=1$ and $\beta=3$. For Str2Str, we use their public codebase to with initial structure guess generated by the ESMFold (Lin et al., 2023) using the script provided in the repository. We use a $T=1000$ diffusion sampling steps for both probability flow and SDE (noise scale is set to $1.0$ for SDE runs) and other parameters as default. For idpGAN, we run the inference pipeline following the Jupyter notebook in their open-source repository with the default inference settings for custom proteins.
@@ A.3 Training details
7:     Compute the negative ELBO loss ${\mathcal{L}}(\theta;{\bm{z}}_{t},t,{\bm{c}})$ based on Eq. 7
@@ A.4 Curation of IDP test set
Algorithm 3 Minibatch Gibbs Sampling with Positional Ranking

### v2

@@ 4.1 Conditional Masked Diffusion Language Model
For conformation generation, we now consider the conditional case of masked diffusion. Given the amino acid types ${\bm{c}}$, our goal is to sample structure tokens through Eq. (2), utilizing a conditional posterior $q({\bm{z}}_{s}|{\bm{z}}_{t},{\bm{z}}_{0};{\bm{c}})$. This posterior can be re-parameterized similarly by incorporating the condition into the backbone model, resulting in $p_{\theta}({\bm{z}}_{s}|{\bm{z}}_{t};{\bm{c}})=q({\bm{z}}_{s}|{\bm{z}}_{t},{\bm{u}}_{\theta}(t,{\bm{z}}_{t},{\bm{c}}))$. To achieve this goal, the reverse process simulated by $p_{\theta}({\bm{z}}_{s}|{\bm{z}}_{t};{\bm{c}})$ must effectively approximate the data distribution $p({\bm{z}}|{\bm{c}})$. A feasible training objective is to optimize the estimation of the conditional ELBO within the continuous-time integral, resulting in the following loss (see Appendix LABEL:app:loss for details):
@@ 4.2 Bidirectional Encoder as denoising network
We now discuss the implementation of the conditional denoising network using bidirectional encoder language models, such as BERT (Devlin, 2018). First, consider the sequential generalization of masked diffusion with a sequence of categorical variables. Let ${\bm{z}}_{t}$ now be a sequence of discrete structure tokens $\left[{\bm{z}}_{t,[1]},{\bm{z}}_{t,[2]},\dots,{\bm{z}}_{t,[L]}\right]$ where ${\bm{z}}_{t,[i]}\in\bar{V},\forall i=1,\dots,L$. Due to the interpolation scheme of Eq. (2), we assume conditional independence and factorize the posterior distribution $p_{\theta}({\bm{z}}_{s}|{\bm{z}}_{t};{\bm{c}})$ across the $L$ output tokens, such that $p_{\theta}({\bm{z}}_{s}|{\bm{z}}_{t},{\bm{c}})=\prod_{i=1}^{L}p_{\theta}({\bm{z}}_{s,[i]}|{\bm{z}}_{t},{\bm{c}})=\prod_{i=1}^{L}q\left({\bm{z}}_{s,[i]}|{\bm{z}}_{t,[i]},{\bm{u}}_{\theta,[i]}(t,{\bm{z}}_{t},{\bm{c}})\right)$ where ${\bm{u}}_{\theta,[i]}(t,{\bm{z}}_{t},{\bm{c}})$ represents the $i$-th output channel of neural network. This implement coincides with the BERT-style transformer architecture and allow us to take advantage of existing protein foundation model, for example ESM3 (Hayes et al., 2024). For a sequence of tokens, the masked $\log$-term in the training objective from Eq. (3) is replaced by the summation: $\sum_{i=1}^{L}\lambda_{M}({\bm{z}}_{t,[i]})\log\langle{\bm{u}}_{\theta,[i]}(t,{\bm{z}}_{t},{\bm{c}}),{\bm{z}}_{0,[i]}\rangle$, with notations the same as defined above.
@@ 5 Experiments
Base settings. We start with the pre-trained dVAE established in Hayes et al. (2024) as the structure tokenizer (frozen). The structure quantization is perform residue-level with a receptive field over local geometric neighborhoods of protein structure. The structure language models as described in Section 3.2 are based off state-of-the-art language models as follows: (1) S-T5 (384M) adopts the T5 (Raffel et al., 2020) architecture with a bidirectional encoder and an autoregressive decoder. (2) S-GPT (961M) models the joint distribution with an uni-directional decoder like GPT2 (Radford et al., 2019). (3) ESM3 (zero shot) is a pre-trained BERT model over multi-modal data and perform zero-shot inference by the iterative decoding. (4) ESMDiff is a fine-tuned variant of (3) on the PDB data using the masked diffusion objective in Eq. (3). For S-T5 and S-GPT, we embed the sequence tokens with the pre-trained ESM3-1.4B encoder to provide model condition. For ESMDiff, two different paradigms: Iterative Decoding (ID) and DDPM are considered. See Appendix B for details.
Table 1: Evaluation results of different conformation generation methods on generating the BPTI conformations. JS-* represents the Jensen-Shannon (JS) divergence between the sampled ensemble and long MD ensemble on pairwise distance (PwD), time-lagged independent components (TIC), and radius of gyration (Rg). The validity as the frequency of clash-free samples in the ensemble is also included. Moreover, the ensemble TM-score and ensemble RMSD are reported with respect to the five kinetic clusters in Shaw et al. (2010). The best results among all methods are bold while the best results among the SLM family are underlined (similarly applied to all tables below).
@@
Baselines. We consider multiple open-source models as evaluation baselines for the protein multiple conformation generation, which are mainly categorized into (1) MSA-based methods that includes MSA-Subsampling (Del Alamo et al., 2022; Bryant & Noé, 2024) and AlphaFlow (Jing et al., 2024a). These methods rely on inference-time retrieval of multiple sequence alignments (MSA) ; (2) Single sequence-based methods: EigenFold (Jing et al., 2023) leverages a harmonic diffusion process conditioned on OmegaFold (Wu et al., 2022) embeddings to generate protein structures, Str2Str (Lu et al., 2024b) simulates a round-trip local diffusion conditioned on input structure to explore hypothetical conformations, and ESMFlow (Jing et al., 2024a) replaces AlphaFlow with ESMFold as the backbone; (3) Specially tailored for intrinsically disordered proteins (IDPs) generation of idpGAN (Janson et al., 2023). Results reported for baselines are obtained by re-running the inference pipeline and based on their open-source codes. The detailed pipeline can be found in Appendix A.
Training data. The training data for structure language models are controlled to contain only PDB entries on or before May 1st, 2020. This cutoff is aligned with previous works (Jing et al., 2023; 2024a; Lu et al., 2024b; Hayes et al., 2024) trained on PDB data to make fair comparison. The training set is further filtered to include all monomeric structures with a max resolution of $5.0\mathrm{\AA}$, length ranging from 10 to 1000, which forms a total size of $|{\mathcal{D}}|=112.4\textrm{k}$ as the training data.
@@ 5.1 Structural dynamics of BPTI
In the first experiment, we evaluate by generating the conformations of protein bovine pancreatic trypsin inhibitor (BPTI). The structural dynamic patterns of BPTI are well acknowledged in Shaw et al. (2010) with 1ms-long MD simulations, based on which five kinetic clusters have been revealed. Similar to Lu et al. (2024b), we report the Jensen Shannon (JS) divergence for distributions of pairwise distance (PwD), time-lagged independent components (TIC), and radius of gyration (Rg); the clash-free validity and the ensemble TM-score and root-mean-square-deviation (RMSD) w.r.t. the kinetic clusters. The benchmark results are shown in Table 5. Following Wang et al. (2024), we also evaluate the best distance of the generated samples to each cluster, as shown in Table 5.1. The RMSD and TM-score are both calculated using the TM-score binary (Zhang & Skolnick, 2004) with structural alignment. Note that the Cluster 3 is a difficult remote folding mode (Wang et al., 2024), yet SLMs achieve a significant improvement by modeling with a smaller matching RMSD.
@@
5.2 Conformation changing pairs
We continue the study on the task of modeling and predicting conformational changes in structural proteins. The authors of Jing et al. (2023) have curated two benchmarking sets of pairing data including (1) 77 pairs of fold-switching proteins (Chakravarty & Porter, 2022); and (2) 90 apo/holo pairs with ligand-induced conformational change (Saldaño et al., 2022) to evaluate the modeling capacity of conformation diversity. Following the setting and evaluation metrics of Jing et al. (2023), we randomly sample a few-shot ensemble with five structures per target and evaluate them based on the correlation metrics of residue flexibility and the ensemble TM-score (Zhang & Skolnick, 2004). The evaluation results for both test sets are shown in Table 5.1. We find that MSA-based methods generally achieve better performance than other model families, which highlights the importance of using MSA for generating stable conformation changing protein targets.
@@ 5.4 Runtime analysis
Figure 5: Runtime profiling for SLMs and other baseline methods. The number of parameters and necessary configurations for each model are also remarked for better reference.
@@
5.5 Case Study: Structural Inpainting of Nanobody
Figure 6: Visualization of a Llama VHH domain. The ESMDiff sampled conformations (in khaki) are superimposed on the NMR structures (Renisio et al., 2002) colored in transparent rainbow.
The position-wise masking and unmasking nature inherently enables ESMDiff to perform the inpainting task effectively. Intuitively, the partially masked input (eg. nanobody framework) can be interpreted as an intermediate state ${\bm{z}}_{s},s>0$ in the masked diffusion, where the unknown positions to be modeled are designated as [MASK] naturally in accordance with the formulation. Then partially reversing allows us to restrict the conformation sampling to the specific sub-regions of the protein, analogous to the concept of “inpainting” for images. Here we name the forward-backward inference as the round-trip masked diffusion inference (see Algorithm LABEL:algo:inpainting) for the inpainting task. Here, we illustrate the conformation inpainting capability through a case study that focuses on generating the complementarity-determining regions (CDRs) for a nanobody example. In specific, we explored the inpainting capabilities of ESMDiff using a nanobody derived from Llama (PDB entry: 1G9E).
As shown in Fig. 6, the loop structures generated by the round-trip diffusion process of ESMDiff resemble the native conformations observed in NMR structures (both $N=20$). The inference is conducted in a very efficient speed by only taking $\sim$5.3 $\pm$ 0.3 seconds on a single NVIDIA A100-SXM4-40GB GPU, highlighting the effectiveness of SLMs in performing structural inpainting and promising potentials in the real-world scenarios such as structural analysis or virtual screening.
@@ 6 Conclusion and Limitations
In this work, we propose a novel conformation generation framework of structure language models (SLMs). The overall inference is divided into two stages: conditional sampling of latent structure tokens and roto-translation invariant structure decoding. We develop and train a variety of conditional language models, especially the masked diffusion-based ESMDiff, which is a fine-tuned variant to enhance the capabilities of ESM3 adapting for conformation generation. Unlike existing methods, SLMs perform amortized distribution learning within an invariant latent space, leading to more efficient inference. By alleviating the need for geometric modeling, SLMs can fully exploit the scalability of modern language model architectures and take advantage of advanced hardware optimizations. Benchmarking results across various conformation generation tasks demonstrate the compelling performance and application potential of SLMs. In summary, the proposed method opens up an intriguing and novel research direction for related communities to explore.
Limitations. The current study presents several limitations worth exploring for future works. Firstly, one can design more advanced dVAE architecture to balance between structure disentanglement and reconstruction fidelity. In addition to the discrete latent space, continuous latent space can also be considered, for example using the latent diffusion models (Rombach et al., 2022). Secondly, it is worthwhile exploring alternative SLM instances specially tailored for the sequence-to-structure translation in consideration of proper inductive biases. Additionally, modeling protein side chains and other relevant modality including ligand and RNA by building an atomic structural auto-encoder is also a very appealing research direction. Lastly, the outstanding performance of MSA-based methods also indicates the potential to build MSA-conditioned structure language models.
Reproducibility Statement
For reproducibility, we provide detailed implementation specifics, including the baseline inference pipeline and training/fine-tuning procedures in Appendix A. The sampling pipelines for ESM3 and ESMDiff are thoroughly described in Appendix B, while Appendix LABEL:app:eval outlines the evaluation methods and their references. The theoretical results mentioned in the main text are derived or proven in Appendix LABEL:app:der. The source training and inference code for structure language models in this study are made publicly available at https://github.com/lujiarui/esmdiff, as we believe fine-tuning foundation protein language models can have broader applicability in various downstream tasks including the protein conformation generation.
Acknowledgments
We thank Can Chen and Chenqing Hua for helpful feedback as well as anonymous reviewers for their constructive suggestion and comments. This project is supported by the Natural Sciences and Engineering Research Council (NSERC) Discovery Grant, the Canada CIFAR AI Chair Program, collaboration grants between Microsoft Research and Mila, Tencent AI Lab Rhino-Bird Gift Fund and a NRC Collaborative R&D Project (AI4D-132). This project was also partially funded by IVADO Fundamental Research Project grant PRF-2019-3583139727. YB acknowledges funding from NRC AI4D, CIFAR and CIFAR AI Chair. The computation resource of this project is supported by Mila, the Digital Research Alliance of Canada, and NRC.
@@ A.1 Baseline evaluations
For baselines, we compare our structure language models against: EigenFold (Jing et al., 2023), Str2Str (Lu et al., 2024b), MSA subsampling (Del Alamo et al., 2022), AlphaFlow & ESMFlow (Jing et al., 2024a), and idpGAN (Janson et al., 2023). For MSA subsampling, we leverage the official repository of Del Alamo et al. (2022) under AlphaFold v2.3.2 with an MSA depth of 32, i.e. 32 maximum extra MSAs and 16 maximum MSA clusters (half the depth value (Del Alamo et al., 2022)). For AlphaFlow and ESMFlow, we use the base PDB model (no distillation) with no resampling. The input MSAs for test proteins are queried via the ColabFold server (Mirdita et al., 2022). All other parameters are set to the recommended values according to the original implementation by the authors (Jing et al., 2024a). For EigenFold, we follow the pipeline in Jing et al. (2023) by using OmegaFold (Wu et al., 2022) to make embeddings for the test set and sampling via the pre-trained diffusion model with default parameters, where $\alpha=1$ and $\beta=3$. For Str2Str, we use their public codebase to with initial structure guess generated by the ESMFold (Lin et al., 2023) using the script provided in the repository. We use a $T=1000$ diffusion sampling steps for both probability flow and SDE (noise scale is set to $1.0$ for SDE runs) and other parameters as default. For idpGAN, we run the inference pipeline following the Jupyter notebook in their open-source repository with the default inference settings for custom proteins.
@@ A.3 Training details
7:     Compute the negative ELBO loss ${\mathcal{L}}(\theta;{\bm{z}}_{t},t,{\bm{c}})$ based on Eq. (3)
@@ A.4 Curation of IDP test set
A.5 Training data of baseline models
We list the training data and the corresponding cutoff for all the mentioned models in our study for reference, as shown in Table S3.
Table S3: Training data and base model of different methods in this study.
Method Base/Embedding model Data source Cutoff date
MSA subs. AlphaFold2 (93M) PDB May 1, 2018
AlphaFlow AlphaFold2 (93M) PDB May 1, 2018
ESMFlow ESMFold (3B) PDB May 1, 2020
Eigenfold OmegaFold (900M) PDB Apr. 30, 2020
Str2Str / PDB Jun, 9, 2023
idpGAN / DisProt50 + MD Jun., 2021
ESMDiff ESM3-open (1.4B) PDB May 1, 2020
Appendix B Inference of Bidirectional SLMs
B.1 Zero-shot conformation generation of ESM3
The sampling method for the BERT model was first introduced by Wang & Cho (2019), which treats the bidirectional encoder as a Markov random field graphical model and employs ID sampling to generate new sequences. In Hayes et al. (2024), the author propose to use iterative decoding to enable simultaneous decoding of multiple positions rather than processing them one at a time. Following the authors (Hayes et al., 2024), we refer to this latter method as the iterative decoding (ID) and take advantages of this sampling strategy for zero-shot conformation generation.
In specific, we leveraged the open-source 1.4B ESM3 protein language model (Hayes et al., 2024) for zero-shot conformation sampling. ESM3 is pre-trained on multi-modal data related to protein and contains several tracks of language model heads for prediction as well as input embeddings, including sequence (amino acid types), structure, secondary structure, SASA and functions. In our main experiment, only the sequence and structure tracks are turned on and the input tokens for other tracks are set to the default value (i.e. [MASK] in each vocabulary). For sampling configurations, we use a sampling schedule of 25 steps and set the sampling temperature to be 1.0 without specially pointing out. During sampling, we also adopted the nucleus (top-p) sampling strategy (Holtzman et al., 2019) to improve the quality of samples with a probability threshold of 0.95. Throughout the experimental Section 5, we use the entropy-ranked ID sampling to obtain the results for ESM3 (described below).
Algorithm 3 Iterative Decoding with Positional Ranking